# Associative Transformer

### v3

@@ 3.4 Information Retrieval with Associative Memory
After writing information into the explicit memory, the learned priors serve as attractors for a novel information retrieval component based on associative memory. The objective is to reconstruct the current input patches towards more globally meaningful representations learned and stored in the explicit memory. In particular, we employ a continuous Hopfield network to iteratively decrease the energy of a patch with respect to the learned priors in the explicit memory, enabling memory retrieval for enhanced patch representations and classification task performance. The proposed architecture is an attractor network in the sense that, in every batch, a pattern converges to one of these attractors derived from the priors stored in the explicit memory.
@@
The learned priors in the explicit memory function as attractors within associative memory of a Hopfield network. Attractors have basins of attraction, and any input that enters an attractor’s basin converges to that specific attractor. Attractors usually have the same dimension as input states. Since we employ low-rank priors $\gamma^{t+1}$, a learnable linear transformation $f_{\text{LT}}(\cdot)$ projects the priors into the same dimension $E$ as the input before utilizing them as attractors.
@@
The upscaled priors $f_{\text{LT}}(\gamma^{t+1})$ are stored within the Hopfield network as various attractors to reconstruct the input state $\Xi^{t}$. In particular, the continuous Hopfield network employs an energy function to enable the evolution of patches into more globally meaningful representations with respect to these attractors. We update each patch representation $\xi^{t}\in\Xi^{t}$ by decreasing its energy $E(\xi^{t})$ within associative memory as follows:
@@ 4 Experiments
In this section, we present the settings and extensive empirical results for image classification tasks, including Triangle, CIFAR, and Oxford Pet, as well as the relational reasoning tasks of Sort-of-CLEVR. We conduct comprehensive ablation studies on AiT and investigate different memory initialization methods. Moreover, we demonstrate the efficacy of the bottleneck attention balance loss by visualizing the emergent prior specialization in the explicit memory.
@@ 4.1 Settings
We evaluate model performance on three different sets of datasets: (1) small classification tasks (Triangle (Goyal et al., 2022), CIFAR10, and CIFAR100 (Krizhevsky & Hinton, 2009)), (2) mid-sized classification tasks (Oxford Pet (Parkhi et al., 2012)), and (3) relational reasoning tasks (Sort-of-CLEVR (Santoro et al., 2017)). We train the model on these datasets from scratch using the training split and evaluate using the test split. A detailed description of the datasets can be found in Appendix A.
@@
We investigate three different sizes of model configurations, i.e., Small, Medium, and Base. The Base variant setting is adapted from Vision Transformer (ViT) using 12 layers, 12 attention heads for each layer, a hidden dimension of 768, and an MLP dimension of 3072. The Medium variant has 6 layers, and the Small variant has 2 layers. In what follows, we use brief notation to indicate the model size, for instance, AiT-B means the “Base” variant of the Associative Transformer. Note that the Global Workspace Layer is a light-weight attention layer with only 0.45M parameters. Further details on the various model sizes of these variants are presented in Section 4.2.
@@
The model was evaluated with a test batch size of one. The explicit memory is only trainable during training and frozen during the evaluation. The hyperparameters were chosen based on a grid search. A batch size of 512 was employed for the CIFAR datasets and the Triangle dataset, 128 for the Pet dataset, and 64 for the Sort-of-CLEVR dataset. We utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.01. A cosine learning rate scheduler was implemented with an initial learning rate of 1e-5, a warm-up phase of 5 (15) epochs within a total of 100 (300) epochs, and a minimum learning rate set to 1e-6. The smoothing factor of the exponentially weighted moving average, the coefficient $\sigma$, and the small value $\epsilon$ in the bottleneck balance loss were set to 0.9, 1e-2, and 1e-10, respectively. The CLS token is removed while the pooled representations of the last dense layer are used since using the CLS token undermines results in vision tasks (Wang et al., 2021; Graham et al., 2021). For AiT, we employed a memory slot size of 32 and a bottleneck attention head size of 8. We used a bottleneck size of 512 for CIFAR and Pet, 64 for Triangle, and 256 for Relational Reasoning. We used 32 memory slots for CIFAR, Triangle, and Relational Reasoning, and 128 slots for Pet. Unless otherwise noted, we trained the model for 100 epochs and reported the mean of three individual experiments. All experiments were based on PyTorch and four A100 GPUs, and the code will be made publicly available. Please refer to Appendix C for the detailed experimental settings and hyperparameters.
@@ 4.2 Classification Tasks
Table 1: Performance comparison in image classification tasks.
Methods CIFAR10 (%) CIFAR100 (%) Triangle (%) Average (%) Model Size (M)
AiT-Base 85.44 60.78 99.59 81.94 91.0
AiT-Medium 84.59 60.58 99.57 81.58 45.9
AiT-Small 83.34 56.30 99.47 79.70 15.8
Coordination (Goyal et al., 2022) 75.31 43.90 91.66 70.29 2.2
Coordination-DH 72.49 51.70 81.78 68.66 16.6
Coordination-D 74.50 40.69 86.28 67.16 2.2
Coordination-H 78.51 48.59 72.53 66.54 8.4
ViT-Base (Dosovitskiy et al., 2020) 83.82 57.92 99.63 80.46 85.7
ViT-Small 79.53 53.19 99.47 77.40 14.9
Perceiver (Jaegle et al., 2021) 82.52 52.64 96.78 77.31 44.9
Set Transformer (Lee et al., 2019) 73.42 40.19 60.31 57.97 2.2
BRIMs (Mittal et al., 2020) 60.10 31.75 58.34 50.06 4.4
Luna (Ma et al., 2021) 47.86 23.38 57.26 42.83 77.6
@@
The experiments on image classification tasks include comparisons to a broad range of methods. We employed the author-recommended hyperparameters to re-implement these methods. In particular, for the Coordination method, we investigated several variants with different model configurations: (1) Coordination consists of 4 layers with parameter sharing among different attention layers, (2) Coordination-D is a deeper model with 8 layers using parameter sharing, (3) Coordination-H is a high memory model with 4 layers that employ individual parameters, and (4) Coordination-DH is a high memory model with 8 layers.
@@
The results in Table 4.2 demonstrate that AiT performance increased when scaling it from Small to Base, while the coordination methods could not scale with the increasing model parameter sizes. AiT also achieved better performance compared to Vision Transformers and the other sparse Transformers including Perceiver, Set Transformer, BRIMs, and Luna. Moreover, AiT relies on significantly fewer layers and parameters while outperforming Vision Transformers. Compared to ViT-Base using 12 layers and 85.7M parameters, AiT-Medium has only 6 layers and 45.9M parameters. Nevertheless, AiT-Medium exhibited an average performance of 81.58%, surpassing ViT-Base’s performance of 80.46%. Image classification experiments on the CIFAR and Triangle datasets demonstrate that AiT outperforms SOTA sparse Transformers and Vision Transformers, requiring a much smaller parameter size to achieve better performance.
@@
Furthermore, we extended the evaluation to a mid-sized dataset of Oxford Pet. We trained the model from scratch for 300 epochs. Figure 2 shows that AiT-Small adds 0.9M parameters to ViT-Small while increasing the performance by 13%. Note that the proposed Global Workspace Layer is lightweight with only 0.45M parameters per attention layer. Although ViT-Base is a much larger model with 85.7M parameters and initially performed better, its accuracy quickly dropped after 50 epochs. In contrast, AiT-Medium with 45.9M parameters demonstrates a stable increase in performance and outperforms the larger ViT-Base model.
Figure 2: Comparison on the Pet dataset. AiT-Small adds 0.9M parameters to ViT-Small while increasing the performance by 13%. AiT-Medium with 45.9M parameters demonstrates a stable increase in performance and outperforms the much larger ViT-Base model.
@@ 4.3 Ablation Study
A comprehensive ablation study aims to acquire insights into the functionalities of different components of AiT. In particular, we conducted the following ablations:
  • Reset Memory: The explicit memory is initialized every epoch.
  • W/O Hopfield: The Hopfield network is replaced with cross-attention that shares the same architecture as the multi-head cross-attention in Figure 1.a. The cross-attention takes the input $\Xi^{t}$ as the query and the upscaled priors $f_{\text{LT}}(\gamma^{t+1})$ as the key and value to compute the output $\hat{\Xi}^{t}=\text{MHA}(\Xi^{t},f_{\text{LT}}(\gamma^{t+1}))$. The rationale behind this ablation is grounded in studies that relied on iterative cross-attentions, such as Set Transformer.
  • W/O Memory: The global workspace layer is removed (equivalent to Vision Transformer).
  • W/O Bottleneck: The top-$k$ bottleneck attention is replaced with dense attention.
@@
  • W/O FF: The feedforward network is removed.
Table 2 demonstrates that the bottleneck plays a significant role in improving performance, where its absence led to an sizable decrease in accuracy. Changes to other components, such as Hopfield networks and explicit memory, while not as impactful, still resulted in degraded accuracy. Consequently, the complete model with all the components achieved the highest accuracy.
We further conducted a comparison in terms of FLOPs for Hopfield network ablations against the baseline AiT method in Table 3. We clarify that the Hopfield network was trained using a single forward-backward iteration, making it lightweight (with less than 0.84% FLOPs increase), especially when compared to the cross attention-based method.
Table 2: Comparison based on an ablation study. The results indicate that combining all the components leads to the highest performance in all the tasks.
Models CIFAR10 CIFAR100 Triangle Average (%)
AiT 83.34 56.30 99.47 79.70
Reset Memory 81.94 55.96 99.46 79.12
W/O Hopfield 81.03 54.96 99.44 78.48
W/O Memory 79.53 53.19 99.47 77.40
W/O Bottleneck 75.40 46.53 93.33 73.75
W/O SA 72.72 47.75 99.46 73.31
W/O FF 69.51 40.89 97.61 69.34
Table 3: FLOPs comparison for Hopfield networks.
Ablation FLOPs
AiT 9.64×$\times 10^{8}$
AiT with Hopfield removed 9.56×$\times 10^{8}$
AiT with cross-attention (Hopfield replaced) 1.19×$\times 10^{9}$
4.4 Parameter Efficiency Compared with Coordination
Figure 3: Model size vs. test accuracy for various configurations within a single attention block. Consolidating all the components in the Coordination block (C+LM+HN+SA), i.e., an AiT block, resulted in the best performance of 71.49% while maintaining a compact model size of 1.0M.
Benefits of incorporating the Global Workspace Layer include its parameter efficiency and better performance. We conducted an extensive comparison of AiT’s parameter efficiency by measuring the model size and corresponding test accuracy on the CIFAR-10 dataset. Specifically, we compared a single AiT block to a single Coordination block that consists of two connected multi-head cross-attention components. We explored various configurations by replacing one or all of the cross-attentions with the proposed components in AiT. In particular, in Figure 3, ‘C’ denotes the default Coordination block, ‘LM’ denotes the ablation replacing the first cross-attention in the Coordination block with the proposed low-rank explicit memory, ‘HN’ denotes the ablation replacing the second cross-attention with the proposed Hopfield network component, and ‘SA’ denotes the ablation using the self-attention.
Figure 3 demonstrates that leveraging the low-rank memory (LM) can improve performance and also reduce the model parameter size. Compared to Coordination (C) which achieved an accuracy of 60.41% with a model size of 2.2M, integrating the low-rank memory (C+LM) significantly improved accuracy while reducing the model size to 1M. Moreover, the Hopfield network (HN) based on the energy function brings improvements (C+SA+HN) or retains the model performance (C+LM+HN+SA) while reducing the parameter size. In addition, HN appeared to be effective only when the LM or SA component was applied. We assume that retrieval within the associative memory of the Hopfield network relies on a diverse set of priors, which is learned by the low-rank memory and further enhanced by the self-attention. Consequently, consolidating all three components in the Coordination block (C+LM+HN+SA), i.e., an AiT block, resulted in the best performance of 71.49% while maintaining a compact model size of 1.0M. The results demonstrate that AiT is an efficient sparse representation learner, surpassing the SOTA Coordination method in terms of both test accuracy and parameter efficiency.
@@ 4.5 Explicit Memory Initialization
Table 4: Different memory initialization approaches. The Gaussian distribution was found to give the best result.
Memory initialization methods Accuracy (%)
Gaussian distribution 83.34
Uniform distribution (Graves et al., 2014) 81.92
Identity distribution (Goyal et al., 2022) 78.56
@@
To initialize the explicit memory, we set each slot with values drawn from a specific distribution, including the Gaussian distribution, the uniform distribution (Graves et al., 2014), and the identity distribution (Goyal et al., 2022). In particular, the Gaussian distribution generates random values with a mean of zero and a variance of one. The uniform distribution (Graves et al., 2014) utilizes an upper bound of $\frac{1}{\sqrt{M+D}}$, where $M$ is the memory slot number, and $D$ is the slot size. The identity distribution (Goyal et al., 2022) assigns ones on the diagonal and zeros elsewhere. Table 4 indicates that the Gaussian distribution resulted in the best performance, which potentially prevents specific memory slots from dominating the learning process, for a more balanced memory update in the early training stages.
4.6 Prior Specialization via Bottleneck Attention Balance Loss
The Bottleneck Attention Balance Loss encourages competition among patches, facilitating the selection of diverse patches through the bottleneck. To quantitatively measure efficacy, we computed a diversity score, i.e., the ratio of distinct patches in the selected patches by all priors in the explicit memory. Figure 8 in Appendix E shows an apparent increase in the diversity of patches as training progresses.

### v4

@@ 3.4 Information retrieval with associative memory
After writing information from the input tokens into the explicit memory, the learned priors serve as attractors for a novel information retrieval component based on associative memory. The aim is to reconstruct the current input tokens towards more globally meaningful representations learned and stored in the explicit memory for enhanced token representations. The proposed architecture is an attractor network where a token converges to one of these attractors derived from the priors stored in the explicit memory, by decreasing its energy with respect to the explicit memory.
@@
The learned priors in the explicit memory function as attractors within associative memory of a Hopfield network. Attractors have basins of attraction, and any input that enters an attractor’s basin converges to that specific attractor. Attractors usually have the same dimension as input states. Since we employ low-rank priors $\gamma^{t+1}\in\mathbb{R}^{M\times D}$, a learnable linear transformation $f_{\text{LT}}(\cdot)$ projects the priors into the same dimension $E$ as the input tokens before utilizing them as attractors.
@@
The upscaled priors $f_{\text{LT}}(\gamma^{t+1})$ are stored within the associative memory of a continuous Hopfield network [21, 22] as various attractors to reconstruct the input tokens $\Xi^{t}$. In particular, we employ the Hopfield energy function to evolve the input tokens into more globally meaningful representations with respect to these attractors. The attractors $f_{\text{LT}}(\gamma^{t+1})$ are shared across instances of the input tokens $\xi^{t}\in\Xi^{t}$. We reconstruct the input tokens using the continuous Hopfield network and the learned attractors in explicit memory as follows:
@@ 4 Experiments
In this section, we present the settings and extensive experiment results for classification and relational reasoning tasks. We compare AiT with a broad range of conventional sparse Transformers in terms of model performance and parameter efficiency. Comprehensive ablation studies demonstrate the effectiveness of the bottleneck attention and associative memory for enhanced model performance.
@@ 4.1 Settings
We evaluate model performance based on three different types of datasets: (1) small classification tasks (Triangle [13], CIFAR10, and CIFAR100 [36]), (2) medium and large-sized classification tasks (Oxford Pet [37] and ImageNet100 [38]), and (3) relational reasoning tasks (Sort-of-CLEVR [14]). We train the model on these datasets from scratch using the training split and evaluate using the test split. A detailed description of the datasets can be found in Appendix A.
@@
We investigate three different sizes of model configurations, i.e., Small, Medium, and Base. The Base variant setting employs 12 layers, 12 attention heads for each layer, a hidden dimension of 768, and an MLP dimension of 3072. The Medium variant has 6 layers, and the Small variant has 2 layers. In what follows, we use brief notation to indicate the model size, for instance, AiT-B means the “Base” variant of the Associative Transformer. Note that the added Global Workspace Layer is light-weight using 0.45M parameters.
@@
The hyperparameters were chosen based on a grid search. A batch size of 512 was employed for the CIFAR datasets and the Triangle dataset, 128 for the Pet dataset, and 64 for the Sort-of-CLEVR dataset. We utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.01. A cosine learning rate scheduler was implemented with an initial learning rate of 1e-5, a warm-up phase of 5 (15) epochs within a total of 100 (300) epochs, and a minimum learning rate set to 1e-6. The smoothing factor of the exponentially weighted moving average, the coefficient $\sigma$, and the small value $\epsilon$ in the bottleneck balance loss were set to 0.1, 1e-2, and 1e-10, respectively. We employed a bottleneck size of 512 for CIFAR and Pet, 64 for Triangle, and 256 for Relational Reasoning. We utilized 32 memory slots for CIFAR, Triangle, and Relational Reasoning, and 128 slots for Pet. Unless otherwise noted, we trained the model for 100 epochs and reported the mean of three individual experiments. All experiments were based on PyTorch and four A100 GPUs, and the code would be made publicly available. The detailed experimental settings and hyperparameters are presented in Appendix C.
@@ 4.2 Classification tasks
Methods CIFAR10 (%) CIFAR100 (%) Triangle (%) Average (%) Size (M) #FLOPs
AiT-Base 85.44 $\pm$ 0.31 60.78 $\pm$ 0.25 99.64 $\pm$ 0.14 81.95 91.0 5.77×$\times 10^{9}$
AiT-Medium 84.59 $\pm$ 0.27 60.58 $\pm$ 0.32 99.57 $\pm$ 0.16 81.58 45.9 2.89×$\times 10^{9}$
AiT-Small 83.34 $\pm$ 0.44 56.30 $\pm$ 0.38 99.47 $\pm$ 0.09 79.70 15.8 9.64×$\times 10^{8}$
Coordination [13] 75.31 $\pm$ 0.72 43.90 $\pm$ 0.21 91.66 $\pm$ 0.56 70.29 2.2 1.46×$\times 10^{8}$
Coordination-DH [13] 72.49 $\pm$ 0.61 51.70 $\pm$ 0.75 81.78 $\pm$ 0.59 68.66 16.6 3.15×$\times 10^{8}$
Coordination-D [13] 74.50 $\pm$ 0.68 40.69 $\pm$ 0.39 86.28 $\pm$ 0.73 67.16 2.2 2.91×$\times 10^{8}$
Coordination-H [13] 78.51 $\pm$ 0.58 48.59 $\pm$ 0.72 72.53 $\pm$ 0.35 66.54 8.4 1.59×$\times 10^{8}$
ViT-Base [2] 83.82 $\pm$ 0.17 57.92 $\pm$ 0.40 99.63 $\pm$ 0.15 80.46 85.7 5.60×$\times 10^{9}$
ViT-Medium [2] 82.41 $\pm$ 0.11 55.78 $\pm$ 0.09 99.62 $\pm$ 0.04 79.27 42.7 2.81×$\times 10^{9}$
ViT-Small [2] 79.53 $\pm$ 0.36 53.19 $\pm$ 0.37 99.47 $\pm$ 0.07 77.40 14.9 9.36×$\times 10^{8}$
Perceiver [26] 82.52 $\pm$ 0.82 52.64 $\pm$ 0.44 96.78 $\pm$ 0.32 77.31 44.9 2.37×$\times 10^{9}$
Set Transformer [23] 73.42 $\pm$ 0.43 40.19 $\pm$ 0.53 60.31 $\pm$ 0.29 57.97 2.2 1.11×$\times 10^{8}$
BRIMs [39] 60.10 $\pm$ 0.50 31.75 $\pm$ 0.28 58.34 $\pm$ 0.43 50.06 4.4 1.43×$\times 10^{8}$
Luna [25] 47.86 $\pm$ 0.53 23.38 $\pm$ 0.06 57.26 $\pm$ 0.19 42.83 77.6 5.08×$\times 10^{9}$
Table 1: Performance comparison in the classification tasks.
Methods Test Accuracy (%) Size (M)
AiT-Medium 36.72 45.9
AiT-Small 33.84 15.8
ViT-Base 34.62 85.7
ViT-Medium 31.72 42.7
ViT-Small 28.16 14.9
Table 2: Performance comparison between AiT and ViT on the ImageNet100 dataset.
@@
The experiments on image classification tasks include comparisons to a broad range of methods. We employed the author-recommended hyperparameters to re-implement these methods. In particular, for the Coordination method [13], we investigated the various model configurations: (1) Coordination consists of 4 layers with parameter sharing among different attention layers, (2) Coordination-D is a deeper model with 8 layers using parameter sharing, (3) Coordination-H is a high memory model with 4 layers that employ individual parameters, and (4) Coordination-DH is a high memory model with 8 layers.
@@
Table 1 demonstrates that AiT performance increased when scaling it from Small to Base, while the Coordination method could not scale with the increasing parameter sizes. AiT achieved better performance compared to Vision Transformers and the other sparse Transformers including Perceiver, Set Transformer, BRIMs, and Luna. Moreover, AiT relies on significantly fewer layers and parameters while outperforming Vision Transformers. Compared to ViT-Base using 12 layers and 85.7M parameters, AiT-Medium has only 6 layers and 45.9M parameters. Nevertheless, AiT-Medium exhibited an average performance of 81.58%, surpassing ViT-Base’s performance of 80.46%. Table 2 shows additional comparison between AiT and ViT on the ImageNet100 dataset. AiT achieved superior performance in the various classification tasks using a much smaller parameter size.
Furthermore, compared to ViT-Base with 5.77×$\times 10^{9}$ FLOPs during training, AiT-Medium utilized only 2.89×$\times 10^{9}$ FLOPs. We compared with the other baseline methods, where the Coordination does not scale well with more model complexity and FLOPs, resulting in degraded performance. In contrast, the proposed Global Workspace Layer is a lightweight architecture, greatly enhancing the classification task performance and parameter efficiency of the conventional sparse Transformers.
Figure 2: Comparison on the Pet dataset. AiT-Medium demonstrated a stable increase in performance outperforming the ViT-Base.
@@
We extended the evaluation to the Oxford Pet dataset [37]. We trained the model from scratch for 300 epochs. Figure 2 demonstrates that AiT-Small with only additional 0.9M parameters (two Global Workspace Layers) increased ViT-Small’s performance by 13%. Although ViT-Base is a much larger model with 85.7M parameters with the optimized model regularizations, its accuracy quickly dropped after 50 epochs. In contrast, AiT-Medium with 45.9M parameters demonstrated a stable increase in performance and outperformed the ViT-Base model.
@@ 4.3 Ablation study
Models CIFAR10 (%) CIFAR100 (%) Triangle (%) Average (%)
AiT 83.34 $\pm$ 0.44 56.30 $\pm$ 0.38 99.47 $\pm$ 0.09 79.70
Reset Memory 81.94 $\pm$ 0.41 55.96 $\pm$ 0.39 99.46 $\pm$ 0.04 79.12
W/O Attention Balance Loss 81.89 $\pm$ 0.39 54.72 $\pm$ 0.33 99.44 $\pm$ 0.08 78.68
W/O Hopfield 81.03 $\pm$ 0.45 54.96 $\pm$ 0.28 99.44 $\pm$ 0.03 78.48
W/O Memory 79.53 $\pm$ 0.36 53.19 $\pm$ 0.37 99.47 $\pm$ 0.07 77.40
W/O Bottleneck 75.40 $\pm$ 0.48 46.53 $\pm$ 0.41 93.33 $\pm$ 0.15 73.75
W/O SA 72.72 $\pm$ 0.57 47.75 $\pm$ 0.31 99.46 $\pm$ 0.04 73.31
Table 3: The ablation study demonstrated that leveraging all the components resulted in the best performance.
Models #FLOPs
AiT 9.64×$\times 10^{8}$
Hopfield network component 8.02×$\times 10^{6}$
W/O Hopfield (replaced with cross-attention) 1.19×$\times 10^{9}$
Table 4: Hopfield network FLOPs.
Initialization Accuracy (%)
Gaussian distribution 83.34 $\pm$ 0.44
Uniform distribution [31] 81.92 $\pm$ 0.17
Identity distribution [13] 78.56 $\pm$ 0.29
Table 5: Memory initialization methods.
A comprehensive ablation study is aimed to acquire insights into the functionalities of different components in AiT. In particular, we conducted the following ablations:
  • Reset Memory: The priors in the explicit memory are initialized every epoch.
  • W/O Hopfield: The Hopfield network is removed and replaced with cross-attention that shares the same architecture as the multi-head cross-attention in Figure 1.a. The cross-attention takes the input $\Xi^{t}$ as the query and the upscaled priors $f_{\text{LT}}(\gamma^{t+1})$ as the key and value to compute the output $\hat{\Xi}^{t}=\text{MHA}(\Xi^{t},f_{\text{LT}}(\gamma^{t+1}))$. The rationale behind this ablation is grounded in studies that relied on iterative cross-attentions, such as Set Transformer.
  • W/O Memory: The global workspace layer is removed.
  • W/O Attention Balance Loss: The bottleneck is implemented without the attention balance loss.
  • W/O Bottleneck: The top-$k$ bottleneck attention is replaced with dense pair-wise attention.
@@
Table 3 demonstrates that the bottleneck plays a significant role in enhancing performance, where its absence led to a sizable decrease in accuracy. The ‘W/O Attention Balance Loss’ablation presents the effectiveness of the proposed bottleneck attention balance loss. Ablating the other components, such as the Hopfield networks and the explicit memory, while not as impactful, still resulted in degraded accuracy. Consequently, the complete model with all the components achieved the best performance. Additionally, for the ‘W/O Hopfield’ablation, Table 4 demonstrates the training time #FLOPs of the Hopfield network compared to the cross-attention component. The Hopfield network is a lightweight component utilizing less than 0.84% FLOPs of the proposed AiT model.
@@ 4.4 Explicit memory initialization
In the explicit memory, we initialized priors with values drawn from a specific distribution, including the Gaussian distribution, the uniform distribution [31], and the identity distribution [13]. In particular, the Gaussian distribution generates random values with a mean of zero and a variance of one. The uniform distribution utilizes an upper bound of $\frac{1}{\sqrt{M+D}}$, where $M$ denotes the number of priors, and $D$ is the dimension of a prior. The identity distribution assigns ones on the diagonal and zeros elsewhere. Table 5 indicates that the Gaussian distribution resulted in the best performance, which potentially prevents specific priors from dominating the learning process, for a more balanced memory update in the early training stages.
4.5 Parameter efficiency
Benefits of incorporating the Global Workspace Layer include enhanced model performance and parameter efficiency. We conducted extensive experiments on AiT’s parameter efficiency by measuring the model size and test accuracy in CIFAR-10. In particular, we compared a single AiT block to a Coordination block that consists of two multi-head cross-attention components. We explored various configurations by replacing the cross-attention components in the Coordination with the proposed low-rank memory and Hopfield network in AiT.
Figure 3: Model size vs. test accuracy for various model configurations. Consolidating all the components in the Coordination block resulted in the best performance of 71.49% maintaining a compact model size of 1.0M.
In Figure 3, ‘C’ denotes the default Coordination block, ‘LM’ denotes replacing the first cross-attention in the Coordination block with the low-rank explicit memory, ‘HN’ denotes replacing the second cross-attention with the Hopfield network, and ‘SA’ denotes the self-attention component. Compared to the Coordination (C) which achieved an accuracy of 60.41% with a model size of 2.2M, integrating the low-rank memory (C+LM) significantly improved accuracy while reducing the model size to 1M. Moreover, the Hopfield network (HN) achieved competitive performance in C+SA+HN and C+LM+HN+SA configurations for enhanced parameter efficiency. In addition, HN appeared to be effective only when the LM or SA component was employed. This is attributed to the retrieval within the associative memory of the Hopfield network relying on learning a diverse set of priors, which are learned based on the low-rank memory and the self-attention components. Removing these components decreases the model’s ability to learn meaningful priors. Consequently, consolidating all the three components in the Coordination block (C+LM+HN+SA), i.e., an AiT block, resulted in the best performance of 71.49% maintaining a compact size of 1.0M. AiT is a sparse representation learner, surpassing the Coordination method in terms of parameter efficiency.
4.6 Prior specialization